the total attendence at a rally was 750. tickets bought before the rally were $2 each,and tickets bought at the door were $2.75 each. How many of ticket were bought at the door if th total receipts were 1706.25
Answer by wuwei96815(245)

Let x=tickets purchased before the rally.
Let 750-x=tickets purchased at the door.

Since we know the total receipts, we can write the following equation:
$2(x) + $2.75(750-x) = $1706.25
2x + 2062.50 - 2.75x = 1706.25
-0.75x = -2062.50 + 1706.25 = -356.25
x = 475
750-x = 275

I hope that you were able to follow the solution.